OpenStudy (anonymous):

What is the solution of log(2x − 6) 256 = 4?

OpenStudy (anonymous):

@mathmale

OpenStudy (abmon98):

Divide both sides by 256 \[\log(2x-6)=4/256\] \[a=\log _{b}(b^a)\] \[\log _{10}(2x-6)=\log _{10(10^4/256}\] They share the same base \[2x-6=10^1/64\]

OpenStudy (anonymous):

This is the question, sorry I might have wrote it wrong @Abmon98

OpenStudy (anonymous):

@Hero

OpenStudy (abmon98):

\[\log_{2x-6} 256=4\] \[\log_{2x-6} 256=\log_{2x-6} (2x-6)^4\]\[256=(2x-6)^4\] \[(256)^1/4=2x-6\] \[4=2x-6\]\[2x=10\]\[x=5\]

OpenStudy (anonymous):

Wow, thank you so much :D

OpenStudy (abmon98):

your welcome :D